### v1

@@ Abstract
Single-molecule localization microscopy generates point clouds corresponding to fluorophore localizations. Spatial cluster identification and analysis of these point clouds are crucial for extracting insights about molecular organization. However, this task becomes challenging in the presence of localization noise, high point density, or complex biological structures. Here, we introduce MIRO (Multimodal Integration through Relational Optimization), an algorithm that uses recurrent graph neural networks to transform the point clouds in order to improve clustering efficiency when applying conventional clustering techniques. We show that MIRO supports simultaneous processing of clusters of different shapes and at multiple scales, demonstrating improved performance across varied datasets. Our comprehensive evaluation demonstrates MIRO’s transformative potential for single-molecule localization applications, showcasing its capability to revolutionize cluster analysis and provide accurate, reliable details of molecular architecture. In addition, MIRO’s robust clustering capabilities hold promise for applications in various fields such as neuroscience, for the analysis of neural connectivity patterns, and environmental science, for studying spatial distributions of ecological data.
@@
Graph neural networks, deep learning, clustering, single-molecule localization microscopy, point clouds, molecular localizationIntroduction
@@
Several algorithms have been specifically proposed for this task [16, 17, 18, 19, 20, 21, 22] and their performance has been recently assessed [23]. Among the methods evaluated in Ref.[23], density-based spatial clustering of applications with noise (DBSCAN) [24], one of the most popular algorithms used for SMLM data, has been shown [23] to be adaptable to diverse clustering conditions and to provide close-to-optimal performance, comparable to those obtained with the topological mode analysis tool (ToMATo)[21] and kernel density estimation (KDE). DBSCAN was also found to be the most robust to multiple blinking. However, DBSCAN’s performance is highly dependent on the choice of its two parameters: the maximum distance between two points for them to be considered as part of the same cluster (); and the minimum number of points that must be within a point’s -neighborhood for that point to be considered a core point and thus form a cluster (minPts). These parameters determine what constitutes a cluster and what constitutes noise. Their choice can significantly affect the resulting clusters, and they require careful dataset-specific settings based on heuristic rules [18, 25] or further analysis [26].
@@
In this paper, we introduce a novel approach to enhance the versatility of clustering algorithms. Our method, MIRO (Multimodal Integration through Relational Optimization), employs a few-shot (or one-shot) geometric deep learning framework based on recurrent graph neural networks (rGNNs) to learn a transformation that squeezes elements of complex point clouds around a common center (Figures 1b–c). To achieve this, MIRO assumes that clusters’ general structure and spatial relationships are preserved within a given dataset and uses relational information to make complex data more suitable for conventional clustering techniques. In this way, MIRO transforms the point clouds so that any method for complete clustering (i.e., assigning every localization to a specific cluster or to the non-clustered group [23]) can achieve enhanced performance, as we demonstrate on a wide range of datasets with varied cluster shape and symmetry. By enhancing the spatial separation between clustered and background localizations, MIRO inherently simplifies the parameters selection for DBSCAN and similar methods. Additionally, the recurrent structure of MIRO inherently supports the simultaneous clustering of structures at different scales, increasing the breadth of information that can be extracted from a single experiment.
@@
Beyond SMLM, MIRO’s robust clustering capabilities hold promise for applications in fields such as neuroscience, where it can help map complex neural networks, providing insights into brain connectivity and function [27], and environmental science, to analyze spatial patterns in ecological data, such as the distribution of species or pollutants, enhancing our understanding of environmental dynamics [28].
Results
MIRO workflow
MIRO uses relational information to transform point clouds to bring together points that belong to the same cluster. It achieves this by using a rGNN, which incorporates several innovative aspects in the architecture, operational mechanisms, and training process, as described here. A detailed description is provided in the Online Methods.
MIRO is built on an rGNN architecture [29]. The input to the neural network is a graph representation of individual molecular localizations derived from SMLM experiments [30]. As shown in Figure 1a, these localizations are obtained from multiple fluorescence images of the same field of view, with each image capturing a sparse number of simultaneously emitting fluorophores. Importantly, fluorophores’ emission is stochastic, therefore a given fluorophore can be detected in multiple frames or not at all. The images are processed to extract the centroid positions of bright features corresponding to molecular localizations. These positions are then drift-corrected and filtered to remove low-quality localizations. Additionally, localizations that are too close together within the same field are discarded, while those that appear in consecutive frames are merged to ensure an accurate representation of distinct molecules.
@@ Results
To strengthen the ability to capture complex spatial relationships, the graph is encoded into a higher-level representation, through learnable linear transformations. The latter serves as the input of a sequence of identical MIRO blocks , as shown in Figure 1c.
@@
The operations of a MIRO block are schematically illustrated in Figure 1d. At each step, the graph is concatenated with a “hidden” graph having the same structure and initialized to zeros. Similar to the hidden state of a recurrent neural network, represents the latent state of the system and characterizes the underlying processes being modeled, capturing relational information between adjacent localizations. Information is propagated to generate an updated hidden graph that is passed to the next step together with the unmodified . In contrast to typical message passing schemes [32, 29, 33], in MIRO, hidden node features only depend on hidden edge features to emphasize the immediate structural context of each node. The hidden node features are further decoded through learnable linear transformations to provide, for each molecular localization, a displacement vector in Cartesian space. These displacements are calculated to minimize a loss function that aims to shift localizations belonging to the same cluster toward a common center, while leaving background localization unaltered.
@@
To ensure a meaningful latent space and prevent vanishing gradients, the loss is calculated at each step [29], as shown in Figure 1c. This approach imposes intermediate corrections to the displacement vectors, helping maintain the clusters’ structural integrity throughout the training process. This method also allows for different steps in the process to have different ground truths, enabling the network to learn and adapt to multiscale features — like the circular clusters () and the ring structures () shown in the example of Figure 1c. Such multiscale training enhances MIRO’s ability to handle varying cluster sizes, shapes, and densities within the same dataset, further improving its robustness and accuracy in clustering complex biological data.
@@
Notably, MIRO’s training can be effectively performed using a single or a few representative clusters. This approach uses the weak conservation of shape and organization within molecular clusters to boost clustering accuracy. By employing a single-shot learning approach combined with a series of augmentations, the algorithm generalizes well across a given scenario, enabling robust performance even when trained on minimal data.
MIRO enhances DBSCAN performance
@@
To demonstrate the benefits of using MIRO, we first applied it to simulated datasets, as illustrated in Figure 2. MIRO is designed as a preprocessing step to enhance the performance of subsequent clustering methods. To assess the performance gains introduced by MIRO, we compared the results of DBSCAN both with and without MIRO preprocessing. We selected DBSCAN for this comparison due to its top performance in benchmark studies [23] and its widespread use in the literature [25].
For DBSCAN without MIRO preprocessing, parameters were selected based on guidelines provided in Ref. [23]. In cases where such guidelines were not available, parameters were optimized based on metric scores, following the procedure described in Ref. [23]. These parameters were consistently applied across all experiments within the same scenario. In contrast, when using MIRO preprocessing, the transformation of clusters into compact, well-defined point clouds increases the spatial separation between clustered and background localizations and thus reduces the influence of minor variations in parameter values on clustering results. For example, MIRO simplifies the selection of the DBSCAN neighboring distance which otherwise requires further analysis of the distance distribution [25] to be set properly. Consequently, parameters for DBSCAN with MIRO preprocessing were selected through visual inspection of the MIRO outputs.
@@
In addition to these metrics, we used cluster-level metrics such as the Jaccard Index for cluster detection (JI), the root mean squared relative error in the number of localizations per cluster (RMSREN), and the root mean squared error in cluster centroid position (RMSEx,y). A summary of these metrics for all scenarios is presented in Table 1.
@@
| Scenario 8 | MIRO | 0.82 ± 0.04 | 0.68 ± 0.06 | 0.85 ± 0.08 | 0.22 ± 0.08 | 2.4 ± 0.3 | 0.84 ± 0.02 | 0.92 ± 0.04 | 0.81 ± 0.03 |
| DBSCAN | 0.81 ± 0.04 | 0.66 ± 0.06 | 0.80 ± 0.11 | 0.3 ± 0.2 | 2.6 ± 0.3 | 0.83 ± 0.03 | 0.90 ± 0.07 | 0.80 ± 0.04 | |
| Scenario 8 blinking | MIRO | 0.85 ± 0.04 | 0.62 ± 0.06 | 0.59 ± 0.08 | 0.3 ± 0.2 | 2.7 ± 0.5 | 0.76 ± 0.03 | 0.88 ± 0.05 | 0.60 ± 0.07 |
| DBSCAN | 0.82 ± 0.05 | 0.59 ± 0.05 | 0.52 ± 0.06 | 0.4 ± 0.2 | 2.8 ± 0.6 | 0.76 ± 0.04 | 0.74 ± 0.08 | 0.64 ± 0.06 | |
| Scenario 9 | MIRO | 0.57 ± 0.05 | 0.71 ± 0.03 | 0.79 ± 0.08 | 0.6 ± 0.4 | 4.2 ± 0.4 | 0.64 ± 0.02 | 0.80 ± 0.07 | 0.50 ± 0.03 |
| DBSCAN | 0.55 ± 0.03 | 0.58 ± 0.04 | 0.72 ± 0.09 | 0.8 ± 0.9 | 4.9 ± 0.8 | 0.62 ± 0.02 | 0.57 ± 0.08 | 0.58 ± 0.02 | |
| Scenario 9 blinking | MIRO | 0.60 ± 0.04 | 0.62 ± 0.03 | 0.62 ± 0.08 | 0.7 ± 0.4 | 4.5 ± 0.7 | 0.64 ± 0.02 | 0.67 ± 0.07 | 0.57 ± 0.03 |
| DBSCAN | 0.59 ± 0.04 | 0.60 ± 0.04 | 0.56 ± 0.09 | 0.9 ± 0.6 | 5.3 ± 0.9 | 0.61 ± 0.03 | 0.56 ± 0.11 | 0.55 ± 0.04 | |
| Scenario 5 blinking | MIRO | 0.59 ± 0.03 | 0.55 ± 0.02 | 0.68 ± 0.03 | 1.0 ± 0.2 | 3.82 ± 0.13 | 0.62 ± 0.02 | 0.64 ± 0.04 | 0.30 ± 0.04 |
| DBSCAN | 0.56 ± 0.03 | 0.54 ± 0.03 | 0.55 ± 0.05 | 1.4 ± 0.2 | 3.91 ± 0.14 | 0.61 ± 0.02 | 0.28 ± 0.05 | 0.43 ± 0.03 | |
| Scenario 6 blinking | MIRO | 0.67 ± 0.04 | 0.65 ± 0.03 | 0.86 ± 0.05 | 0.35 ± 0.09 | 4.4 ± 0.5 | 0.74 ± 0.02 | 0.83 ± 0.05 | 0.65 ± 0.04 |
| DBSCAN | 0.65 ± 0.03 | 0.61 ± 0.03 | 0.72 ± 0.09 | 0.6 ± 0.2 | 4.8 ± 0.5 | 0.70 ± 0.02 | 0.65 ± 0.06 | 0.65 ± 0.03 | |
| C-shaped | MIRO | 0.95 ± 0.02 | 0.97 ± 0.01 | 0.999 ± 0.005 | 0.06 ± 0.03 | 0.27 ± 0.06 | 0.97 ± 0.01 | 0.95 ± 0.02 | 0.94 ± 0.02 |
@@
In our evaluation, we observed that MIRO significantly refines DBSCAN’s performance across various scenarios. We first focused on assessing MIRO on datasets from the benchmark study [23], characterized by different cluster density, size, and shape (Figures 2a–c). For instance, in Scenario 8 (small symmetrical clusters with two different densities), while the scatter and box plots in Figure 2a show that MIRO only slightly improves the performance of DBSCAN, this improvement is statistically significant. This scenario represents a case where the performance of DBSCAN is close-to-optimal, therefore, it is not surprising that MIRO only makes a small difference. Specifically, MIRO achieves a small to medium effect size (Cohen’s ) for both ARI† and IoU, with this improvement being most pronounced in cluster-level metrics such as JIc and RMSREN.
@@
As expected, the advantage of using MIRO becomes more evident in more challenging conditions. In Scenario 8 with blinking (Figure 2b), the increased number of localizations due to molecular overcounting introduces more heterogeneity into the data, but MIRO effectively mitigates this effect and significantly improves DBSCAN’s performance (Cohen’s for ARI†). A similar improvement is observed in Scenario 9 (Table 1), which features clusters of different sizes. MIRO further demonstrates its capability to handle additional complexities when, in addition to blinking, the number of clusters is increased, as in Scenario 5 (Figure 2c).
@@
To further highlight MIRO’s ability in managing complex cluster geometries, we evaluated its performance under three additional conditions. First, we examined Scenario 6 with blinking, which includes elliptically-shaped clusters. Additionally, we simulated data with C-shaped and ring-shaped clusters (Figure 2d-f). In these scenarios, MIRO produces a marked enhancement in DBSCAN’s performance by consistently transforming elongated and non-convex shapes into well-defined, compact clusters for the further application of DBSCAN.
Simultaneous clustering and classification of different shapes

### v2

@@ Abstract
Single-molecule localization microscopy generates point clouds corresponding to fluorophore localizations. Spatial cluster identification and analysis of these point clouds are crucial for extracting insights about molecular organization. However, this task becomes challenging in the presence of localization noise, high point density, or complex biological structures. Here, we introduce MIRO (Multifunctional Integration through Relational Optimization), an algorithm that uses recurrent graph neural networks to transform the point clouds in order to improve clustering efficiency when applying conventional clustering techniques. We show that MIRO supports simultaneous processing of clusters of different shapes and at multiple scales, demonstrating improved performance across varied datasets. Our comprehensive evaluation demonstrates MIRO’s transformative potential for single-molecule localization applications, showcasing its capability to revolutionize cluster analysis and provide accurate, reliable details of molecular architecture. In addition, MIRO’s robust clustering capabilities hold promise for applications in various fields such as neuroscience, for the analysis of neural connectivity patterns, and environmental science, for studying spatial distributions of ecological data.
@@
Graph neural networks, deep learning, clustering, single-molecule localization microscopy, point clouds, molecular localization1 Comments
This is the author’s version of the article published in Nature Communications under a Creative Commons Attribution 4.0 International License (CC BY 4.0). The final published version is available at https://doi.org/10.1038/s41467-025-65557-7.
2 Introduction
@@
Several algorithms have been specifically proposed for this task [16, 17, 18, 19, 20, 21, 22] and their performance has been recently assessed [23]. Among the methods evaluated in Ref.[23], density-based spatial clustering of applications with noise (DBSCAN) [24], one of the most popular algorithms used for SMLM data, has been shown to be adaptable to diverse clustering conditions and to provide close-to-optimal performance, comparable to those obtained with the topological mode analysis tool (ToMATo)[21] and kernel density estimation (KDE). DBSCAN was also found to be the most robust to multiple blinking [23]. More recently, DBSCAN has been shown to achieve significantly higher scores than HDBSCAN [25] and OPTICS [26] across different cluster types [27]. However, DBSCAN’s performance is highly dependent on the choice of its two parameters: the maximum distance between two points for them to be considered as part of the same cluster (); and the minimum number of points that must be within a point’s -neighborhood for that point to be considered a core point and thus form a cluster (minPts). These parameters determine what constitutes a cluster and what constitutes noise. Their choice can significantly affect the resulting clusters, and they require careful dataset-specific settings based on heuristic rules [18, 28] or further analysis [29, 27].
@@
In this paper, we introduce a novel supervised approach to enhance the versatility of clustering algorithms. Our method, MIRO (Multifunctional Integration through Relational Optimization), employs a few-shot (or one-shot) geometric deep learning framework based on recurrent graph neural networks (rGNNs) to learn a transformation that squeezes elements of complex point clouds around a common center (Figure 1b–c). To achieve this, MIRO assumes that clusters’ general structure and spatial relationships are preserved within a given dataset and uses relational information to make complex data more suitable for conventional clustering techniques. In this way, MIRO transforms the point clouds so that methods for complete clustering (i.e., assigning every localization to a specific cluster or to the non-clustered group [23]) can achieve enhanced performance, as we demonstrate for DBSCAN on a wide range of datasets with varied cluster shape and symmetry. By enhancing the spatial separation between localizations in adjacent clusters, as well as between clustered and background localizations, MIRO inherently simplifies the parameter selection for DBSCAN and similar methods. Additionally, the recurrent structure of MIRO facilitates a multifunctional representation framework, enabling the simultaneous handling of heterogeneous analysis tasks, such as multiscale clustering, clustering of differently shaped structures, and node-level classification. Unlike traditional pipelines that treat these tasks separately or require manual tuning, MIRO learns a flexible representation that supports diverse objectives in a unified and scalable manner. This multifunctional capability significantly expands the range of biologically relevant insights that can be extracted from a single experiment.
@@
3 Results
3.1 MIRO workflow
MIRO uses relational information to transform point clouds to bring together points that belong to the same cluster. It achieves this by using a rGNN, which incorporates several innovative aspects in the architecture, operational mechanisms, and training process, as described here. A detailed description is provided in the Methods.
MIRO is built on an rGNN architecture [30]. The input to the neural network is a graph representation of individual molecular localizations derived from SMLM experiments [31]. As shown in Figure 1a, these localizations are obtained from multiple fluorescence images of the same field of view (FOV), with each image capturing a sparse number of simultaneously emitting fluorophores. Importantly, fluorophores’ emission is stochastic, therefore a given fluorophore can be detected in multiple frames or not at all. The images are processed to extract the centroid positions of bright features corresponding to molecular localizations. These positions are then drift-corrected and filtered to remove low-quality localizations. Additionally, localizations that are too close together within the same field are discarded, while those that appear in consecutive frames are merged to ensure an accurate representation of distinct molecules.
@@ 3.1 MIRO workflow
To strengthen the ability to capture complex spatial relationships, the graph is encoded into a higher-level representation through a learnable dense layer followed by ReLU activation. The latter serves as the input of a sequence of recurrent steps, were the same MIRO block is repeatedly applied, as shown in Figure 1c. We emphasize that MIRO uses a single-layer architecture. As a result, increasing the number of recurrent steps does not affect the number of learnable parameters. The number of recurrent steps defines the size of the receptive field and therefore needs to be adapted to the density of the point cloud and the complexity of the clustering problem, as discussed in Number of recurrent steps: influence on performance and oversmoothing.
@@
The operations of a MIRO block are schematically illustrated in Figure 1d. At each recurrent step, the graph is concatenated with a “hidden” graph having the same structure and with node and edge features initialized to zeros. Similar to the hidden state of a recurrent neural network, represents the hidden state of the system and characterizes the underlying processes being modeled, capturing relational information between nearby localizations. Information is propagated to generate an updated hidden graph that is passed to the next step together with the unmodified . In contrast to typical message passing schemes [33, 30, 34], MIRO omits the concatenation of node features with aggregated messages. Instead, hidden node features are updated solely based on hidden edge features (i.e., the messages) to emphasize the immediate structural context of each node. The hidden node features are further decoded through a learnable dense layer to provide, for each molecular localization, a displacement vector in Cartesian space. These displacements are calculated to minimize a loss function (see MIRO loss function) that aims to shift localizations belonging to the same cluster toward a common center, while leaving background localization unaltered.
@@
To ensure a meaningful hidden representation and prevent vanishing gradients, at each iteration within an epoch, the loss is further averaged across all recurrent steps [30], as schematically shown in Figure 1c. This approach imposes intermediate corrections to the displacement vectors, helping to maintain the clusters’ structural integrity throughout the recurrent steps. This method also allows for different steps in the process to have different ground truths, enabling the network to learn and adapt to multiscale features — like the circular clusters () and the ring structures () shown in the example of Figure 1c. Such multiscale training enhances MIRO’s ability to handle varying cluster sizes, shapes, and densities within the same dataset, further improving its robustness and accuracy in clustering complex biological data.
@@
Notably, MIRO’s training can be effectively performed using a single or a few representative clusters (see MIRO training and augmentations). This approach uses the weak conservation of shape and organization within molecular clusters to boost clustering accuracy. By employing a series of augmentations, the algorithm learns to generalize across a given scenario, enabling robust performance even when trained on minimal data.
3.2 MIRO enhances DBSCAN performance
@@
To demonstrate the benefits of using MIRO, we first applied it to simulated datasets, as illustrated in Figure 2. MIRO is designed as a preprocessing step to enhance the performance of subsequent clustering methods. To assess the performance gains introduced by MIRO, we compared the results of DBSCAN both with and without MIRO preprocessing. We selected DBSCAN for this comparison due to its top performance in benchmark studies [23, 27] and its widespread use in the literature [28].
For the benchmark datasets, we used the DBSCAN parameters provided in Ref. [23]. For MIRO-preprocessed data and other datasets, clustering parameters were optimized using an automated procedure based on the Optuna Python library [35], guided by metric-based performance scores. These parameters were consistently applied across all experiments within the same scenario and are summarized in Supplementary Table 1 and Supplementary Table 2. Please refer to DBSCAN parameter selection for a discussion of the parameter choice criteria.
@@
In addition to these metrics, we used cluster-level metrics such as the Jaccard Index for cluster detection (JIc), the root mean squared relative error in the number of localizations per cluster (RMSREN), and the root mean squared error in cluster centroid position (RMSEx,y).
@@
| Scenario 8 | MIRO | 0.83 ± 0.04 | 0.72 ± 0.06 | 0.88 ± 0.08 | 0.16 ± 0.06 | 2.3 ± 0.3 | 0.85 ± 0.03 | 0.92 ± 0.04 | 0.83 ± 0.03 |
| DBSCAN | 0.81 ± 0.04 | 0.65 ± 0.06 | 0.80 ± 0.11 | 0.3 ± 0.2 | 2.6 ± 0.3 | 0.83 ± 0.03 | 0.90 ± 0.07 | 0.80 ± 0.04 | |
| Scenario 8 blinking | MIRO | 0.84 ± 0.04 | 0.62 ± 0.06 | 0.57 ± 0.07 | 0.28 ± 0.13 | 2.6 ± 0.4 | 0.77 ± 0.05 | 0.81 ± 0.08 | 0.65 ± 0.08 |
| DBSCAN | 0.82 ± 0.05 | 0.59 ± 0.05 | 0.52 ± 0.05 | 0.4 ± 0.2 | 2.9 ± 0.6 | 0.76 ± 0.04 | 0.74 ± 0.08 | 0.64 ± 0.06 | |
| Scenario 5 blinking | MIRO | 0.59 ± 0.02 | 0.56 ± 0.02 | 0.70 ± 0.04 | 0.78 ± 0.13 | 3.71 ± 0.14 | 0.62 ± 0.02 | 0.50 ± 0.05 | 0.37 ± 0.03 |
| DBSCAN | 0.56 ± 0.03 | 0.54 ± 0.03 | 0.56 ± 0.05 | 1.3 ± 0.2 | 3.92 ± 0.15 | 0.60 ± 0.02 | 0.28 ± 0.05 | 0.42 ± 0.03 | |
| Scenario 6 blinking | MIRO | 0.66 ± 0.04 | 0.66 ± 0.03 | 0.89 ± 0.06 | 0.40 ± 0.12 | 4.3 ± 0.5 | 0.74 ± 0.02 | 0.84 ± 0.05 | 0.65 ± 0.03 |
| DBSCAN | 0.65 ± 0.03 | 0.61 ± 0.03 | 0.72 ± 0.09 | 0.6 ± 0.2 | 4.8 ± 0.5 | 0.70 ± 0.02 | 0.65 ± 0.06 | 0.66 ± 0.03 | |
| C-shaped | MIRO | 0.95 ± 0.02 | 0.968 ± 0.011 | 0.999 ± 0.005 | 0.06 ± 0.03 | 0.27 ± 0.06 | 0.967 ± 0.011 | 0.95 ± 0.02 | 0.94 ± 0.02 |
@@
In our evaluation, MIRO consistently enhances the performance of DBSCAN across all tested scenarios, as shown in Supplementary Table 3 and Supplementary Table 4. While these results are based on few-shot training (three FOVs, comprising 60 to 300 clusters), we also demonstrate that comparable performance can be achieved with single-shot training, as shown in Supplementary Table 5 for representative scenarios.
@@
First, we discuss MIRO’s performance on selected datasets from the benchmark study [23], characterized by different cluster density, size, and shape (Table 1 and Figure 2a–c). For instance, in Scenario 8 (small symmetrical clusters with two different densities), while the scatter and box plots in Figure 2a show that MIRO only slightly improves the performance of DBSCAN, this improvement is statistically significant. This scenario represents a case where the performance of DBSCAN is close-to-optimal, therefore, it is not surprising that MIRO only makes a small difference. Specifically, MIRO achieves a medium effect size for ARI† (Cohen’s ; paired one-sided Wilcoxon test , , ) and a large effect size for IoU (Cohen’s ; , , ), with this improvement being most pronounced in cluster-level metrics such as JIc (Cohen’s ; , , ) and RMSREN (Cohen’s ; , , ).
@@
As expected, the advantage of using MIRO becomes more evident in more challenging conditions. In Scenario 8 with blinking (Figure 2b), the increased number of localizations due to molecular overcounting introduces more heterogeneity into the data, but MIRO effectively mitigates this effect and significantly improves DBSCAN’s performance (Cohen’s ; , , for ARI†). MIRO further demonstrates its capability to handle additional complexities when, in addition to blinking, the number of clusters is increased, as in Scenario 5 (Cohen’s ; , , for ARI†, Figure 2c).
@@
To further highlight MIRO’s ability in managing complex cluster geometries, we evaluated its performance under three additional conditions. First, we examined Scenario 6 with blinking, which includes elliptically-shaped clusters. Additionally, we simulated data with C-shaped and ring-shaped clusters (Figure 2d–f). In these scenarios, MIRO produces a marked enhancement in DBSCAN’s performance by consistently transforming elongated and non-convex shapes into well-defined, compact clusters for the further application of DBSCAN.
We further demonstrate that MIRO outperforms recent supervised methods not included in the benchmark, such as an implementation of the GNN-based framework proposed in Ref. [22] (MAGIK-S), as shown in Supplementary Table 6. Details of the implementation are provided in Comparison with a supervised graph-based clustering framework.
3.3 Simultaneous clustering and classification of different shapes